# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v2

@@ Abstract.
Large language models (LLMs) increasingly play an important role in a wide range of information processing and management tasks. Many of these tasks are performed in large batches or even offline, and the performance indictor for which is throughput. These tasks usually show the characteristic of prefix sharing, where different prompt input can partially show the common prefix. However, the existing LLM inference engines tend to optimize the streaming requests and show limitations of supporting the large batched tasks with the prefix sharing characteristic. The existing solutions use the LRU-based cache to reuse the KV context of common prefix between requests. The KV context that are about to be reused may prematurely evicted with the implicit cache management. Besides, the streaming oriented systems do not leverage the request-batch information and can not mix the decoding tokens with the prefill chunks to the best for the batched scenarios, and thus fails to saturate the GPU. We propose BatchLLM to address the above problems. BatchLLM explicitly identifies the common prefixes globally. The requests sharing the same prefix will be scheduled together to reuse the KV context the best. BatchLLM reorders the requests and schedules the requests with larger ratio of decoding first to better mix the decoding tokens with the latter prefill chunks, and applies memory-centric token batching to enlarge the token-batch sizes, which helps to increase the GPU utilization. Finally, BatchLLM optimizes the prefix-shared Attention kernel with horizontal fusion to reduce tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × on a set of microbenchmarks and a typical industry workload under different hardware environments.
1. Introduction
@@
The modern information processing and management tasks are starting to leverage large language models (LLMs) to achieve better results, such as the search engine (Wang et al., 2023; Xiong et al., 2024), advertisement (Meguellati et al., 2024), and recommender systems (Zhao et al., 2024; Kim et al., 2024). Due to the huge amount of data to be processed, many of these LLM tasks are performed in large batches or even offline (e.g., offline ranking), for which the most important performance indicator is throughput. Besides, different input (prompts) of these tasks can partially share the same prefix (e.g., the same web document in the search engine tasks). Given the high cost of LLM inference, how to serve these tasks efficiently is a critical problem for the information processing systems.
@@
The recent LLM inference engines (Kwon et al., 2023; team, 2024) can flexibly support the execution of LLM inference tasks, especially with blocked KV memory management (Kwon et al., 2023) and per-iteration token batching (Yu et al., 2022). However, the design of existing inference engines tend to optimize streaming online services and show limitations in throughput-oriented large-batch processing. 1) Lack of global prefix sharing. The state-of-the-art (SOTA) LLM inference engine (Kwon et al., 2023) supports to cache the KV context with LRU policy to enable prefix sharing between prompts. However, from a global perspective of large batches, an LRU-based cache may prematurely evict KV contexts that are about to be reused and retain unused KV contexts for a long time, causing unnecessary recalculation. 2) Suboptimal token batching for throughput-oriented and prefix-shared tasks. The current LLM inference engines schedule different requests independently and do not cluster the requests with common prefix together to schedule. It can extend the lifetime of the common prefix and thus increase the memory usage, and may prematurely evict KV contexts that are about to be reused as discussed above. Besides, they usually schedule the requests in the first-come-first-serve or similar order to guarantee the fairness and latency of the streaming requests. The requests with larger ratio of decoding steps may be scheduled too late to be able to mix with the prefill chunks to increase the hardware utilization. They apply the mix of prefill and decoding tokens together according to the threshold of token number and request number. This limits the number of decoding tokens that can be batched and keeps the GPU111 This paper illustrates the technique with the GPU architecture. But the basic principle can be applied to other architectures. from saturating for the iterations dominated by decoding tokens. 3) Room for improvement in Attention. The SOTA prefix-shared Attention optimization can be further optimized with horizontal kernel fusion of the computation on different KV chunks, which is effective to reduce the tail effect and kernel launch overhead.
@@
In this work, we introduce BatchLLM, the holistic optimization for the throughput-oriented large-batch LLM inference with global prefix preprocessing, prefix-aware and throughput-oriented token batching and Attention kernel optimization. The basic insight is based on the fact that, the prompt characteristics are known ahead before processing the large batch. Compared to the implicit LRU cache that cannot retain the KV context for reuse accurately, the common prefixes between prompts can be identified explicitly and can be reused explicitly. Besides, the length distribution of the prompts is known ahead and the main performance target is throughput, which allows the inference engine to reorder the requests and form larger token-batches222 This paper uses the term token-batch to represent a batch of tokens to be processed for an iteration of continuous batching (background at Sec.2.3). It differs from the concept of large batch of all the prompt to be processed of a job. .
Building upon these insights, this paper proposes a set of optimizations. It builds the global prefix tree on the large batch ahead-of-time to enable the explicit prefix sharing and avoid the problem of premature eviction of KV contexts. To simplify the design and reduce the overhead of token batching and Attention kernel on multi-level prefix, it simplifies the multi-level prefix into a single level one while retaining the similar KV reusing ratio with Dynamic Programming algorithm on the tree. It schedules the requests sharing the common prefix together to enable them reusing the KV context. To form large token-batches at each iteration to increase GPU utilization, it reorders the requests ahead-of-time according to the length of prompt and the decoding (estimated). The request with larger ratio of decoding length to prompt length will be scheduled first so that the decoding tokens will be processed earlier and can be mixed with the latter longer prefill chunks. It also batches the tokens according to the KV memory usage rather than only depending on request number threshold, to form large batches for decoding tokens. Finally, it applies horizontal fusion of Attention calculation on different KV chunks to reduce the overhead of tail effect and kernel launch of the basic prefix-shared Attention implementation.
We have implemented BatchLLM on top of vLLM (Kwon et al., 2023) and conducted extensive evaluations using a set of microbenchmarks of different datasets and models and an industry workload on NVIDIA and AMD GPUs. BatchLLM achieves end-to-end performance improvement of 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × than vLLM and SGLang for the microbenchmarks and an industry workload under different hardware environments. The performance gain comes from BatchLLM’s ability to reuse the common prefix explicit in the global view, form the token-batches with higher compute-intensity, and the more efficient prefix-shared Attention.
In summary, this paper makes the following contributions:
@@
2. LLM Inference Background
2.1. Performance Factors of LLM Inference
The auto-regressive LLM inference usually consists of two phases: the prefill phase to process the input prompt, and the decoding phase to generate the output. The prefill phase is usually compute-intensive and bounded by the massive computation. This is because the input tokens are all known ahead, and the computation of these tokens can be performed in parallel. In the decoding phase, the output tokens are generated one by one and cannot be performed in parallel333 In this paper, we mainly focus on the basic decoding scheme and will not discuss the parallel decoding technique (Cai et al., 2024).. As a result, the decoding phase can be easily bounded by the memory access of the weight due to the memory wall problem (Xia et al., 2023, 2024). Enlarging the batch size can help to push the decoding computation to compute-bound, but the batch size may be limited by the GPU memory capacity. In real practice, the batching of the decoding can be suboptimal due to either the lack of memory space or the inefficient token batching, leading to low hardware utilization. Attention can be an important performance factor when the sequence is long. According to its algorithm pattern, the basic Attention computation in decoding phase is the matrix-vector multiplication between the (batched) vector $Q$ and matrix $KV$ containing all history context, as tokens from different requests have different KV context. Due to this characteristic, enlarging the batch size usually can not increase the compute utilization of the GPU as it does not increase the compute-intensity.
2.2. KV Cache and Prefix Sharing
Based on the nature of Attention calculation (Vaswani et al., 2017), the processing of each token (in both prefill and decoding phase) will generate the corresponding KV context. The computation of latter tokens consume the KV context of earlier tokens to perform Attention calculation. During the processing of a request (i.e., a prompt), the KV context is usually cached in memory for each of the processed token (called KV cache) so that the latter tokens can reuse the history KV context in memory for Attention calculation. The KV cache can consume a large amount of GPU memory, which can limit the batch size of LLM inference and thus hurt the GPU hardware utilization of the linear layers consist of matrix multiplication (MatMul) calculation. The KV cache is usually managed in blocked memory to reduce fragmentation and enable KV sharing (Kwon et al., 2023), the logical continuous KV tensors are mapped to the physical discrete memory blocks with the block table.
When different prompts share the same prefix, the common KV context ($K_{prefix}$, $V_{prefix}$) can be computed once and reused by these prompts. The Attention computation between $Q$ and $K/V$ can be performed through the partial Attention on the prefix and the distinct prompt respectively, followed by the Online Softmax (Milakov and Gimelshein, 2018) to reduce the partial results:
(1) $\begin{split}O_{prefix}&=partial\_attention(Q,K_{prefix},V_{prefix})\\ O_{distinct}&=partial\_attention(Q,K_{distinct},V_{distinct})\\ O&=online\_softmax(O_{prefix},O_{distinct})\end{split}$
The existing LLM inference engines make use of compact prefix tree (i.e., Radix tree) and blocked KV memory to enable KV sharing of the common prefix between different prompts (Kwon et al., 2023; Zheng et al., 2023). Due to that different prompts can partially share the same KV context of the prefix, the partial Attention calculation on the prefix ($K_{prefix}$, $V_{prefix}$) can be transformed from the matrix-vector mulplication into MatMul to increase the compute intensity (Ye et al., 2024b; Zhu et al., 2024; Juravsky et al., 2024; Ye et al., 2024a). The existing works perform the calculations in Eq.1 in three separated GPU kernels. If there are multiple levels of prefix, each of the extra prefix will introduce two more kernels for the partial Attention and reduction respectively (Ye et al., 2024a).
Besides the saved computation of the common prefix, by storing the KV context of the common prefix with a single copy, it can help to decrease the memory pressure caused by KV memory and potentially increase the allowed batch size.
2.3. Per-iteration Token Batching
Due to the dynamic output length of LLM execution, batching the tasks at the granularity of request results in a waste of the GPU resource, as some requests may finish earlier. Instead, the current LLM inference engines batch the tasks at the granularity of tokens. Specifically, it forms the token-batch at each iteration and allows to add new tokens into the batch dynamically (Yu et al., 2022), which is also called continuous batching. During the execution, the linear operators of different tokens will be concatenated together into a larger linear operator for execution at each iteration.
Recent works (Holmes et al., 2024; Agrawal et al., 2023) split the long prefill into chunks and mix the prefill chunks with decoding tokens into the same token-batch to conquer the memory-wall problem of pure decoding batches, called SplitFuse or chunked-prefill. By inserting the prefill chunks into the decoding batches, the token number of the batch is enlarged effectively without introducing a lot of memory consumption. As a result, the compute-intensity of the token-batches are increased and the hardware utilization becomes better. Mixing the decoding tokens with prefill chunks is an effective way to increase the throughput, but can potentially increase the latency of the decoding tasks, thus preventing existing LLM inference engines from enabling chunked-prefill by default.
3. Opportunities and Insight
3.1. Emerging Industry Scenarios
@@
LLM has been increasingly used in a broad range of industry tasks with its in-context learning and reasoning ability (Wang et al., 2023; Xiong et al., 2024; Meguellati et al., 2024; Zhao et al., 2024; Kim et al., 2024; Chen et al., 2023; Fernandez et al., 2023). Due to the large traffic of these tasks and the high cost of LLM inference, many of these tasks are processed in very large batch (or even offline) whose performance metric is mainly the throughput rather than latency. Take the snippet generation task (Wang et al., 2023; Xiong et al., 2024) in search engine as an example, which is to generate the snippet of the web document (web content) in the search result page according to both document and user query. An example prompt in a previous work (Wang et al., 2023) is: generate a short snippet based on the given Document to answer the given Query. Document: <…> Query: <…>. Due to that it is hard to meet the SLO when performing the massive LLM inferences online, a common practice is to run the snippet generation task offline for the high-frequency web documents and queries, and retrieve the offline results during online serving when the corresponding document-query pair occurs.
@@
Many of the LLM-based information processing and management tasks show the characteristic of the shared prefix between the prompts of different requests, which comes from the nature that different tasks can perform on the same content (e.g., web document). Take the search engine task as an example, it may extract different information from the same web document so that the document can be a shared prefix in the prompt. Specifically, as for the snippet generation task, given that it is to extract the snippet of the document-query pair, a document can be the common prefix for different queries.
@@
The shared prefix can be managed in multiple levels. For example, the first level prefix can be the global system information and instructions, and the second level can be the web document. However, with the development of LLM and the using of task-specific model fine-tuning, the prefix of system information and instructions becomes shorter and shorter, and the length of the prefix will finally be dominated by the web document itself. On the one hand, the LLM models are incorporating more and more reasoning abilities into themselves (like OpenAI o1 (OpenAI, 2024a)) and will not rely on complex prompts. On the other hand, the task-specific fine-tuning can help to incorporate the complex prompts into the model weight. As a result, the prefix sharing of the long context rather than the instructions can be the most important for many industry tasks. This motivates the first-level prefix enlargement in Sec.4.2.
3.2. The New Optimization Demands
@@
As discussed in Sec.2, the existing works have proposed solutions to support prefix sharing and per-iteration token batching to support LLM inference. However, the state-of-the-art solutions still lack specialized optimization for the increasingly import large batched scenarios described in Sec.3.1, showing three main limitations:
@@
1. The implicit prefix caching cannot lead to the optimal KV reuse for the large batched inference. The existing LLM serving systems (Kwon et al., 2023; Zheng et al., 2023) use the prefix tree to maintain the KV blocks. The same prefix (identified by runtime hashing) can be mapped to the same KV blocks to avoid repeated computation. It uses the LRU policy to evict the blocks from the block table. As for a large batch of inputs to be processed, it does not keep the prefix in memory in the global view in the large batch, but only according to the history input. As a result, the shareable KV blocks can be easily evicted prematurely. We have evaluated vLLM’s prefix sharing with a typical industry workload. The optimal token saving ratio by prefix reusing is 58.1% by analyzing the input dataset. Note that the saving ratio is calculated by:
(2) $R_{saving}=(1-\frac{N_{processed\_prefill\_tokens}}{N_{logical\_prefill\_tokens}})\times 100$
@@
$N_{processed\_prefill\_tokens}$ is the number of prefill tokens processed after prefix sharing, and $N_{logical\_prefill\_tokens}$ is the original prefill token number of all the requests. The vLLM’s implicit prefix caching only achieves 35.8% (refer to Sec.6.3.1). There is a demand to enable better prefix sharing for the large batched tasks, especially considering that all the prefill characteristics in the batch are known ahead.
@@
Figure 1. The effect of processing order of requests with chunked-prefill enabled. Given the three requests with different prefill/decoding length characteristics, the naive token batching in the coming order of the requests has worse token mixing of decoding and prefill chunks.
@@
2. The online serving oriented scheduling and token batching can be suboptimal for the prefix-shared and throughput oriented tasks. The current LLM inference systems schedule the tasks at the granularity of request. They do not analyze the prefix sharing characteristics of the whole large batch nor schedule the requests sharing common prefix together. This not only can evict the shareable KV context prematurely as discussed above, but also extends the lifetime of the shareable KV memory, exacerbating the problem of large KV memory consumption. Besides, these systems design tends to support online services. They form the token-batches at each iteration with the constraint of requests’ arriving order, which can result in suboptimal token-batches. Take the example in Fig.1, the naive scheduling in the order of request’s coming cannot mix the decoding of request 3 with the prefill chunks of other requests (chunked-prefill background in Sec.2.3). Instead, by scheduling request 3 earlier, its decoding can be mixed with the prefill chunks of other requests. Another problem is that, the current systems use the token and request number in the token-batch as the threshold to limit the batching, which limits to batch more tokens together in the decoding dominated token-batches and prevents better utilizing the GPU even when the memory is enough. Fig.2 shows the number of tokens at each iteration for a typical industry task (Sec.6.2.2) performing with vLLM’s best configuration. It shows that there are "valleys"444 This paper uses valley to express the token-batches that has smaller number of tokens, appearing as the valley in the timeline of token numbers. at many iterations where the number of tokens are not large enough to saturate the GPU. There is a demand to reschedule the tasks to make the requests sharing common prefix closer, and enable better mix of decoding tokens with prefill chunks. There is also a room to enlarge the size of token-batches dominated by the decoding tokens.
@@
Figure 2. The token number in the batch processed at each iteration for an industry task with vLLM’s chunked-prefill. It has "valleys" for many iterations.
3. There is still room for Attention optimization of prefix sharing. As described in Sec.2.2, the existing works (Ye et al., 2024b; Zhu et al., 2024; Juravsky et al., 2024; Ye et al., 2024a) support prefix-shared Attention by computing the Attention on different chunks of KV, performing MatMul on the common prefix and the basic matrix-vector multiplication on the non-shared chunk. But the computation on different chunks are in different kernels. On the one hand, the separated kernels increase the tail effect of GPU kernel. On the other hand, the multiple kernels introduce extra launching overhead.
3.3. Insight

### v3

@@ Abstract
Large language models (LLMs) increasingly play an important role in a wide range of information processing and management tasks in industry. Many of these tasks are performed in large batches or even offline, and the performance indicator for which is throughput. These tasks usually show the characteristic of prefix sharing, where different prompt input can partially show the common prefix. However, the existing LLM inference engines tend to optimize the streaming requests and show limitations of supporting the large batched tasks with the prefix sharing characteristic. The existing solutions use the LRU-based cache to reuse the KV context of common prefix between requests. The KV context that are about to be reused may be prematurely evicted with the implicit cache management. Besides, the streaming oriented systems do not leverage the request-batch information and can not mix the decoding tokens with the prefill chunks to the best for the batched scenarios, and thus fails to saturate the GPU. We propose BatchLLM to address the above problems. BatchLLM explicitly identifies the common prefixes globally. The requests sharing the same prefix will be scheduled together to reuse the KV context the best. BatchLLM reorders the requests and schedules the requests with larger ratio of decoding first to better mix the decoding tokens with the latter prefill chunks, and applies memory-centric token batching to enlarge the token-batch sizes, which helps to increase the GPU utilization. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by $1.3\times$ to $10.8\times$ on a set of microbenchmarks and a typical industry workload under different hardware environments. Code is available at https://github.com/microsoft/MixLLM/tree/batchllm_vllm_064.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
@@
The modern information processing and management tasks are starting to leverage large language models (LLMs) to achieve better results, such as the search engine Wang et al. (2023); Xiong et al. (2024), advertisement Meguellati et al. (2024), and recommender systems Zhao et al. (2024); Kim et al. (2024). Due to the huge amount of data to be processed, many of these LLM tasks are performed in large batches or even offline (e.g., offline ranking), for which the most important performance indicator is throughput. Besides, different input (prompts) of these tasks can partially share the same prefix (e.g., the same web document in the search engine tasks). Given the high cost of LLM inference, how to serve these tasks efficiently is a critical problem for the information processing systems.
@@
The recent LLM inference engines Kwon et al. (2023); Zheng et al. (2023) can flexibly support the execution of LLM inference tasks, especially with blocked KV memory management Kwon et al. (2023) and per-iteration token batching Yu et al. (2022). However, the design of existing inference engines tend to optimize streaming online services and show limitations in throughput-oriented large-batch processing. 1) Lack of global prefix sharing. The state-of-the-art (SOTA) LLM inference engines support to cache the KV context with LRU policy to enable prefix sharing between prompts. However, from a global perspective of large batches, an LRU-based cache may prematurely evict KV contexts that are about to be reused and retain unused KV contexts for a long time, causing unnecessary recalculation. 2) Suboptimal token batching for throughput-oriented and prefix-shared tasks. The current LLM inference engines schedule different requests independently and do not cluster the requests with common prefix together to schedule. It can extend the lifetime of the common prefix and thus increase the memory usage, and may prematurely evict KV contexts that are about to be reused as discussed above. Besides, they usually schedule the requests in the first-come-first-serve or similar order to guarantee the fairness and latency of the streaming requests. The requests with larger ratio of decoding steps may be scheduled too late to be able to mix with the prefill chunks to increase the hardware utilization. They apply the mix of prefill and decoding tokens together according to the threshold of token number and request number. This limits the number of decoding tokens that can be batched and keeps the GPUs from saturating for the iterations dominated by decoding tokens.
@@
In this work, we introduce BatchLLM, the holistic optimization for the throughput-oriented large-batch LLM inference with global prefix preprocessing, and prefix-aware and throughput-oriented token batching. The basic insight is based on the fact that, the prompt characteristics are known ahead before processing the large batch. Compared to the implicit LRU cache that cannot retain the KV context for reuse accurately, the common prefixes between prompts can be identified explicitly and can be reused explicitly. Besides, the length distribution of the prompts is known ahead and the main performance target is throughput, which allows the inference engine to reorder the requests and form larger token-batches [^1] This paper uses the term token-batch to represent a batch of tokens to be processed for an iteration of continuous batching Yu et al. (2022). It differs from the concept of large batch of all the prompt to be processed in a job. . It makes the following contributions:
@@
2 Opportunities
2.1 Emerging Industry Scenarios
@@
LLM has been increasingly used in a broad range of industry tasks with its in-context learning and reasoning ability Wang et al. (2023); Xiong et al. (2024); Meguellati et al. (2024); Zhao et al. (2024); Kim et al. (2024); Chen et al. (2023); Fernandez et al. (2023). Due to the large traffic of these tasks and the high cost of LLM inference, many of these tasks are processed in very large batch (or even offline) whose performance metric is mainly the throughput rather than latency. Take the snippet generation task Wang et al. (2023); Xiong et al. (2024) in search engine as an example, which is to generate the snippet of the web document (web content) in the search result page according to both document and user query. An example prompt in industry Wang et al. (2023) is: generate a short snippet based on the given Document to answer the given Query. Document: <…> Query: <…>. Due to that it is hard to meet the SLO when performing the massive LLM inferences online, a common practice is to run the snippet generation task offline for the high-frequency web documents and queries, and retrieve the offline results during online serving when the corresponding document-query pair occurs.
@@
Many of the LLM-based information processing and management tasks show the characteristic of the shared prefix between the prompts of different requests, which comes from the nature that different tasks can perform on the same content (e.g., web document). Take the search engine task as an example, it may extract different information from the same web document so that the document can be a shared prefix in the prompt. Specifically, as for the snippet generation task, given that it is to extract the snippet of the document-query pair, a document can be the common prefix for different queries. There are many other industry tasks that fall into the offline processing mode with a common prefix. The very important offline ranking tasks for search and recommendation can use the prompt of "please evaluate the relevance of the document for the query, Query: <…>, Document: <…>", where each query corresponds to many documents and can be the common prefix. The Ads title rewrite task in industry can use the prompt of "given the Query and the Landing Page info, please refine the old title ‘xxx’ to get a better ad title. Landing Page: <…>, Query: <…>", where the landing page can be the common prefix and the computation is usually throughput critical. The offline labeling and scoring tasks in industry can have long system prompt and few shot examples, which can also be the common prefix in the prompt.
@@
The shared prefix can be managed in multiple levels. For example, the first level prefix can be the global system information and instructions, and the second level can be the web document. However, with the development of LLM and the using of task-specific model fine-tuning, the prefix of instructions becomes shorter and shorter: On the one hand, the LLM models are incorporating more and more reasoning abilities into themselves (like OpenAI o1 OpenAI (Cited 2024a)) and will not rely on complex prompts. On the other hand, the task-specific fine-tuning can help to incorporate the complex prompts into the model weight. As a result, the prefix sharing of the long context rather than the instructions can be the most important for many industry tasks. This motivates the first-level prefix enlargement in Sec.3.2.
2.2 The New Optimization Demands
@@
Even though prefix sharing and per-iteration token batching are functionally supported in previous works, they lack specialized optimization for the increasingly import large batched scenarios (Sec.2.1), showing two main limitations:
@@
1. The implicit prefix caching cannot achieve the optimal KV reuse for the large batched inference. The industrial LLM serving systems Kwon et al. (2023); Zheng et al. (2023) use the prefix tree to maintain the KV blocks. The same prefix (identified by runtime hashing) can be mapped to the same KV blocks to avoid repeated computation. It uses the LRU policy to evict the blocks from the block table. As for a large batch of inputs to be processed, it does not keep the prefix in memory in the global view in the large batch, but only according to the history input. As a result, the shareable KV blocks can be easily evicted prematurely. We have evaluated vLLM’s prefix sharing with a typical industry workload. The optimal token saving ratio by prefix reusing is 58.1% by analyzing the input dataset. Note that the saving ratio is calculated by:
$R_{saving}=(1-\frac{N_{processed\_pre\_tok}}{N_{logical\_pre\_tok}})\times 100$ (1)
@@
$N_{processed\_pre\_tok}$ is the number of prefill tokens processed after prefix sharing, and $N_{logical\_pre\_tok}$ is the original prefill token number of all the requests. The vLLM’s implicit prefix caching only achieves 35.8% (Sec.4.3.1). There is a need for better prefix sharing for the large batched tasks.
@@
Figure 1: The effect of processing order of requests with chunked-prefill enabled. Given the three requests with different prefill/decoding length characteristics, the naive token batching in the coming order of the requests has worse token mixing of decoding and prefill chunks.
@@
2. The online serving oriented scheduling and token batching can be suboptimal for the prefix-shared and throughput oriented tasks. The current LLM inference systems schedule the tasks at the granularity of request. They do not analyze the prefix sharing characteristics of the whole large batch nor schedule the requests sharing common prefix together. This not only can evict the shareable KV context prematurely as discussed above, but also extends the lifetime of the shareable KV memory, exacerbating the problem of large KV memory consumption. Besides, these systems design tends to support online services. They form the token-batches at each iteration with the constraint of requests’ arriving order, which can result in suboptimal token-batches. Take the example in Fig.1, the naive scheduling in the order of request’s coming cannot mix the decoding of request 3 with the prefill chunks Holmes et al. (2024); Agrawal et al. (2023) of other requests. Instead, by scheduling request 3 earlier, its decoding can be mixed with the prefill chunks of other requests. Another problem is that, the current systems use the token and request number in the token-batch as the threshold to limit the batching, which limits to batch more tokens together in the decoding dominated token-batches and prevents better utilizing the GPU even when the memory is enough. Fig.2 shows the number of tokens at each iteration for a typical industry task (Sec.4.2.2) performing with vLLM’s best configuration. It shows that there are "valleys" at many iterations where the number of tokens are not large enough to saturate the GPU. There is a demand to reschedule the tasks to make the requests sharing common prefix closer, and enable better mix of decoding tokens with prefill chunks. There is also a room to enlarge the size of token-batches dominated by the decoding tokens.
@@
Figure 2: The token number in the batch processed at each iteration for an industry task with vLLM’s chunked-prefill. It has "valleys" for many iterations.
3 Design Methodology
3.1 Overview